OpenStudy (elisuzsmith):

solve the system by elimination: -2x +2y +3z=0, -2x-y+z=-3, 2x +3y+3z=5

OpenStudy (anonymous):

let -2x +2y +3z=0 be your equation 1 -2x-y+z=-3 be your equation 2 2x +3y+3z=5 be your equation 3 solve eqn 4 by subtracting eqn 1and eqn 2 solve eqn 5 by add eqn 1 and eqn 3 tell me what you got..for eqn 4 and 5